Apples cost $0.60 and oranges cost $0.40 each. Jack bought 5 more apples than oranges, and he paid a total of $10.00. How many apples did he buy?
A - 12
B - 13
C - 11
D - 5
E - 24

Answer by macston(5194)

A=apples; R=oranges=A-5
$0.60A+$0.40R=$10.00 Substitute for R.
$0.60A+$0.40(A-5)=$10.00
$0.60A+$0.40A-$2.00=$10.00 Add $2.00 to each side.
$1.00A=$12.00 Divide each side by $1.00.
A=12 ANSWER: Jack bought 12 apples. Choice A is correct.
CHECK:
R=A-5=12-5=7 Jack bought 7 oranges
$0.60A+$0.40R=$10.00
$0.60(12)+$0.40(7)=$10.00
$7.20+$2.80=$10.00
$10.00=$10.00